OpenStudy (anonymous):

Ruby red paint costs $0.50 per square inch. A ball is painted ruby red and costs $100.00 to paint. What is the diameter of the ball? Round to the nearest tenth of an inch

OpenStudy (math&ing001):

Circumference of the sphere = 2*pi*r (square inch) $100.00 for 2*pi*r (square inch) $0.50 for 1 (square inch) By the rule of 3: 2*pi*r = 100/0.50 Find r, then the diameter is 2r.